Please help with my problem # 19. Write the answers in set-builder notation.
5 - 9x > = 19 + 5x
Thank you.

Answer by sdmmadam@yahoo.com(530)

5 - 9x > = 19 + 5x ----(1)
5-19 > = 5x+9x (grouping like terms,changing sign while changing side)
-14 > = 14x
dividing by 14
(division by a positive quantity does not alter the inequality and hence here greater than remains greater than)
-1 > = x
That is x < = -1
x belongs to the open close interval -infinity to -1
That is x (belongs to) (-infinity,-1]
Note: Please supply the symbol at the right place namely the belongs to symbol and the infinity symbol in your note book.
I don't know how to use them in this answer box.